OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 25? x squared over 25 ±square root of the quantity x plus 25 ±5 times the square root of x 1 over quantity x squared minus 25

OpenStudy (austinl):

\(f(x)=x^2-25\) The first thing you have to do is put it in terms of x and y. \(y=x^2-25\) Then you flip the x and y. \(x=y^2-25\) Then you solve for y. Then replace y with f(x). I am relatively sure that this is how you do it.

OpenStudy (jdoe0001):

\(\bf f(x) ={\color{red}{ y}}= {\color{blue}{ x}}^2 - 25\qquad inverse\implies {\color{blue}{ x}}= {\color{red}{ y}}^2 - 25\) solve for "y" notice that to get the inverse "relation", we simply swap about the variables

OpenStudy (austinl):

replace y with \(f^{-1}(x)\)

OpenStudy (anonymous):

thanks

OpenStudy (austinl):

No problemo! Have a great day.